Two blocks are resting on the floor. Block A is 3.2 kg. Block B is 4.7 kg. Both blocks need to be lifted up onto a benchtop surface that is 75 cm above the floor. How much more work will it take to lift Block B on to the benchtop than to lift Block A?
1
Expert's answer
2019-12-17T09:59:54-0500

The work to lift up block of mass "M"onto a benchtop surface that is "H" above the floor is "A=M\cdot g\cdot H" where "g=9.8 \frac{m}{s^2}" is acceleration of gravity. The difference between two work will be "\delta A= M_2\cdot gH - M_1\cdot gH = (M_2 - M_1)gH"

"\delta A=(4.7 - 3.2)kg\cdot 9.8\frac{m}{s^2}\cdot 0.75 m=11.0 J"

Answer: The work of lifting block B on the benchtop surface is more than when lifting block A by 11.0 J.